OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? y = 3x + 5 y = 3x + 3 y = 3x - 17 y = 3x - 5

OpenStudy (anonymous):

y=3x+3

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (anonymous):

yea sure no problem

OpenStudy (anonymous):

Which relation does not represent a function? (-9, 3), (-2, 1), (1, 3), (5, -6) (2, 0), (4, 8), (7, 5), (1, 5) (-6, 2), (2, 5), (-4, 0), (2, 6) (-4, 6), (-1, 1), (0, 4), (4, 7)

OpenStudy (anonymous):

the second one doesn't represent a function. A function is only One to One.

OpenStudy (anonymous):

it's the second one.

OpenStudy (anonymous):

@agutie I have already said this!!

OpenStudy (anonymous):

Oh wait I see, it has a number in there twice.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes this is correct!

OpenStudy (anonymous):

What is the slope between the two points (4, 14) and (2, 4)? 1/5 5 -1/5 -5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Okay that's what I thought.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which equation is not equivalent to the formula r = st? s=r/t t=r/s s=t/r r = ts I think it's the second one.

OpenStudy (anonymous):

s=t/r

OpenStudy (anonymous):

How? I don't understand these & I'm prepping for an exam.

OpenStudy (anonymous):

if you try to get on variable by itself you do the divsion. and check the answers that are given to you and see if they match or not.

OpenStudy (anonymous):

Ohh! I see now. At least you explain it clearer.

OpenStudy (anonymous):

ok. its not that difficult if you understand it very well.

OpenStudy (anonymous):

If f(x) = x2 + 3, find f(5). 13 22 28 4 I got 28 this is how: f(x)=x^2+3 f(5)=5^2+3 f(5)=25+3 f(5)=28

OpenStudy (anonymous):

@Sceenie yes.!

OpenStudy (anonymous):

Awesome! :D

OpenStudy (anonymous):

:D you get it now.!

OpenStudy (anonymous):

Solve for x: 3x - 2 > 5x + 10 x > 6 x < 6 x < -6 x > -6 I got x < -6 3x-2>5x+10 +2 +2 ---------- 3x>5x+12 -5x -5x ---------- -2x>12 -- -- -2 -2 x < -6

OpenStudy (anonymous):

Yes.!! :D

OpenStudy (anonymous):

Can you explain this one a little bit? The set of ordered pairs below defines a relation. In each ordered pair (x, y), x represents the daily wages of a plumber, and y represents the daily wages of her assistant. {(144, 72), (171, 84), (198, 84), (171, 96), (144, 72)} What is the difference between the values of the least and greatest elements in the domain of this relation? 24 54 72 126

OpenStudy (anonymous):

I think it's 54, cause if I'm understanding it right, I just have to subtract the numbers in the ordered pair.

OpenStudy (anonymous):

54

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

Simplify 4(2u - 3). 8u - 12 8u + 12 8u - 3 6u - 7 I think it's : 8u-3

OpenStudy (anonymous):

8u-12

OpenStudy (anonymous):

you have to distribute the 4 to the whole equation

OpenStudy (anonymous):

Wait lol i forgot to distribute the other side. xD

OpenStudy (anonymous):

yes cx

OpenStudy (anonymous):

Simplify the following expression. 7x - 4y + 5z - 2x - 10y + z 5x - 14y + 6z 9x - 14y + 6z 5x - 6y + 6z 9x - 6y + 6z I think it's :5x-14y+6z 7x - 4y + 5z - 2x - 10y + z 7x-2x-4y-10y+5z+z rx-14y+6z

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

x-6 --- = 5 4 I got x = 26

OpenStudy (anonymous):

yes. sorry. yes that's right